OpenStudy (anonymous):

convert each equation to standard form by completing the square on x and y. Then graph the ellipse and give the location of its foci. 4x^2+y^2+16x-6y-39=0 4x^2+25y^2-24x+100y+36=0

OpenStudy (anonymous):

you need steps or answer?

OpenStudy (anonymous):

first one is \[4(x+2)^2+(y-3)^2=64\] or \[\frac{x+2)^2}{1}+\frac{(y-3)^2}{4}=16\] and that put it in standard form for an ellipse

OpenStudy (anonymous):

thank you